OpenStudy (anonymous):

b) The function f(x)=x^2-7x+1 has a tangent with the slope 3. Calculate the equation for the tangent.

myininaya (myininaya):

The tangent at what point?

myininaya (myininaya):

Yep that could work. So have you found that point yet @z17 ?

OpenStudy (anonymous):

It doesn't say. Let me check again.

myininaya (myininaya):

No I was asking for the point and @Loser66 was telling you how to get it.

myininaya (myininaya):

If we didn't know the slope, how would you go about finding?

OpenStudy (anonymous):

I have the slope of 3?

myininaya (myininaya):

I know, but if we didn't.

myininaya (myininaya):

What does the derivative tell us?

OpenStudy (anonymous):

Whether it's increasing or decreasing?

myininaya (myininaya):

Right and doesn't it also tell you the slope?

OpenStudy (anonymous):

Yes?

myininaya (myininaya):

So find the derivative of your function

OpenStudy (anonymous):

Is it x^2?

myininaya (myininaya):

Your function is x^2-7x+1 I'm asking you to find the derivative of it so we can get the slope.

myininaya (myininaya):

What is the derivative of x^2?

myininaya (myininaya):

Do you know the power rule?

OpenStudy (anonymous):

Yes, it just takes me time to calculate it.

myininaya (myininaya):

The power rule says the derivative of x^n is nx^(n-1)

OpenStudy (anonymous):

I'm so confused to be honest

myininaya (myininaya):

So you don't know the power rule then?

myininaya (myininaya):

Do you know the formal definition of a derivative?

OpenStudy (anonymous):

dy/dx ?

myininaya (myininaya):

Yes what is the formal definition?

OpenStudy (anonymous):

Don't remember -.-

myininaya (myininaya):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h} \text{ formal definition } \]

myininaya (myininaya):

or you could use: \[f'(a)=\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}\]

OpenStudy (anonymous):

So what do I do in order to calculate the equation?

myininaya (myininaya):

I'm still trying to get you to find the derivative of x^2-7x+1 This is step one. We haven't got passed it yet.

OpenStudy (anonymous):

and I do that by using the power rule right?

myininaya (myininaya):

You could but it didn't appear you knew it.

myininaya (myininaya):

And then there is also constant multiple rule and constant rule.

myininaya (myininaya):

So that is why I was going a step backwards to the actual definitions.

myininaya (myininaya):

Since you didn't know the power rule.

OpenStudy (anonymous):

How do I apply the power rule to find the 1st derivate?

myininaya (myininaya):

\[\text{ power rule } (x^n)'=nx^{n-1} \]

myininaya (myininaya):

Use this rule to find the derivative of x^2

myininaya (myininaya):

what is n?

myininaya (myininaya):

When is x^2 equal to x^n?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

I need to study

myininaya (myininaya):

so you do know that x^2 equals x^2

myininaya (myininaya):

so if i asked you when does x^2 equal x^n you would say when n equals ?

OpenStudy (anonymous):

x^2?

myininaya (myininaya):

and what value of n are you stating?

myininaya (myininaya):

\[x^2=x^n => n=? \]

OpenStudy (anonymous):

What do mean by value of n? I'm not used to studying math in english.

myininaya (myininaya):